# Supersymmetry and Localization on Three-Dimensional Orbifolds

### v1

@@ Abstract
We consider three-dimensional ${\mathcal{N}}=2$ supersymmetric field theories defined on general complex-valued backgrounds of Euclidean new minimal supergravity admitting two Killing spinors of opposite R-charges. We focus on compact spaces admitting orbifold singularities, including $\mathbb{\Sigma}\times S^{1}$, where $\mathbb{\Sigma}$ is a spindle, as well as branched and squashed lens spaces. In each of these supersymmetric backgrounds we compute the corresponding partition function, thus obtaining novel observables for three-dimensional supersymmetric gauge theories. In particular, our results cover a wide class of circle bundles over a spindle. For theories compactified on $\mathbb{\Sigma}\times S^{1}$ we discuss both the twist and anti-twist and show that the partition functions on these backgrounds may be expressed in terms of a single formula, which we refer to as the spindle index, unifying and generalising the refined superconformal and topologically twisted indices, in the limits wherein the orbifold singularities are removed. We test our new index by means of non-perturbative dualities. The main results of this paper were announced in [1].
@@ 1 Introduction
The investigation of supersymmetric quantum field theories (SQFTs) stands as a longstanding pursuit in theoretical physics as it represents an invaluable avenue for exploring non-trivial phenomena, particularly in the regime of strong interactions, where perturbation theory fails. From a physics standpoint, the exact computations of SQFTs’ partition functions on curved backgrounds, made possible through localization techniques [2], have produced profound insights on non-perturbative aspects of gauge theories, including the discovery of dualities among seemingly completely different models. From a mathematical point of view, such exact partition functions may serve as generating functions for a number of interesting invariants characterizing the manifolds where the theory is defined (for a recent review see e.g. [3] and references therein).
@@
Partition functions of SQFTs defined on the direct product of a circle and a compact manifold play a special role as they correspond to supersymmetric indices, which may be interpreted in terms of “counting” of various quantities. The study of indices has revealed unforeseen correlations across different areas of physics and mathematics, establishing connections between SQFTs, the geometry of the moduli space in the context of string theory, as well as in purely mathematical fields such as symplectic and algebraic geometry and representation theory, e.g. Gromov-Witten invariants and characters of infinite-dimensional algebras. Recent advancements in localization techniques have yielded significant breakthroughs, facilitating explicit computations of supersymmetric indices as well as other partition functions corresponding to manifolds with the different topology and dimension.
@@
Supersymmetric partition functions play also an important role in the AdS/CFT correspondence, where they provide precision checks of dualities between gravitational backgrounds and holographically dual field theories. The most striking instance of this aspect is the use of supersymmetric indices to perform a precise counting of black hole microstates, thereby offering a glimpse into the fundamental tenets of quantum gravity. In particular, through the study of large-$N$ limits, supersymmetric indices have been used to obtain a microscopic derivation of the entropy of supersymmetric black holes [4, 5, 6, 7]. Several other works have studied different asymptotic limits and determined sub-leading corrections, matching these to higher-derivative or quantum corrections in supergravity [8, 9, 10, 11]. Another set of precision tests in holography involves matching the large-$N$ limit of the partition of function of three-dimensional ${\cal N}=2$ on manifold with $S^{3}$ topology [12, 13, 14] to various supergravity solutions [15, 16, 17, 18, 19].
@@
In the past two years, starting with [20], a substantial body of literature has demonstrated that supergravity solutions displaying orbifold singularities should not be discarded. On the contrary, such solutions are expected to play a crucial role as non-trivial saddle points in the quantum gravity path integral, or in its semi-classical limit. In this extended landscape of solutions supersymmetry can be realized in novel fashions, raising the intriguing question of how typical these supergravity backgrounds are. One of the simplest and perhaps most interesting set-ups is that four-dimensional supersymmetric [^1] See also [21] for a discussion of supersymmetry of the solutions of [22, 23]. and accelerating black holes [22, 23], in which the horizon is a spindle $\mathbb{\Sigma}$ [24], where $\mathbb{\Sigma}=\mathbb{WCP}^{1}_{{\left[{{n_{+}},{n_{-}}}\right]}}$ is the two-dimensional weighted projective space whose topology is that of a two-sphere with conical singularities $\mathbb{C}/\mathbb{Z}_{n_{\pm}}$ at the poles. The conformal boundary of these geometries comprises a metric on $\mathbb{\Sigma}\times\mathbb{R}_{t}$ together with a gauge field with “anti-twisted” flux on $\mathbb{\Sigma}$; therefore, via holography, their microstates are expected to be captured by the supersymmetric partition function of the dual ${\cal N}=2$ three-dimensional field theories defined on a rigid supersymmetric background on $\mathbb{\Sigma}\times S^{1}$ [25]. In this work we will discuss in detail the calculation of this new observable, that we named spindle index in [1], where the main results of the present paper were anticipated.
@@
More generally, the recent supergravity constructions involving orbifolds raised a number of related questions concerning superconformal field theories (SCFTs). In would be interesting to understand systematically the data that define a SCFT at an orbifold point and what are the additional properties that one can hope to detect, on such backgrounds. For SCFTs in even space-time dimensions some progress has been made by studying the ’t Hooft anomalies of the theories compactified on orbifolds, extending anomaly polynomial techniques to spaces with conical singularities [20, 26, 27]. This approach is somewhat insensitive to the detailed description of the microscopic fields at the orbifold singularities, suggesting that a similar point of view may be fruitful for studying field theories on orbifolds also in odd space-time dimensions. In particular, we expect that orbifold supersymmetric indices and other partition functions of field theories should unveil novel features of SQFTs (in odd as well as in even space-time dimension), beyond those that are immediately relevant for applications in holography.
[…]
@@
In Section 2 we characterise a very general class of complex local rigidly supersymmetric backgrounds. We will show that these comprise all the known supersymmetric backgrounds as well as several generalizations, that we discuss in detail. In particular, we describe general backgrounds on $\mathbb{\Sigma}\times S^{1}$, with either twist or anti-twist for the background R-symmetry gauge field. The conformal boundary of the accelerating black hole of [24] as well as its complex deformation [25] are shown to fit into our general canonical form. In Section 3 we compute the full supersymmetric partition function of three-dimensional ${\cal N}=2$ Chern-Simons-matter theories on $\mathbb{\Sigma}\times S^{1}$. The one-loop determinants are obtained employing both the method of unpaired eigenvalues as well as the equivariant orbifold index theorem. We also compute various classical terms and discuss subtleties related to the breaking of gauge invariance and its restoration through the introduction of effecting terms. Our final result for the spindle index is tested with some simple field theory dualities, that hold in a form modified by the presence of the orbifold singularities. We consider both anti-twist and twist and show that the final expression can be written in a unified form, giving the spindle index [1]. We also include flavour fugacities as well as several classical terms, that were not discussed in [1]. As further applications of our formalism, in Section 4 we compute the full supersymmetric partition function of three-dimensional ${\cal N}=2$ Chern-Simons-matter theories on a branched and squashed lens space, again finding a result that unifies and generalises previous constructions regarding non-trivial circle (orbi-)bundles over two-dimensional (orbifold) Riemann surfaces. In particular, our partition function includes the squashed lens space one [31] and the branched sphere one [32]. Features of this observable and its possible applications will be discussed further in Section 6. In Section 5 we show how to recover the one-loop determinants relevant for the spindle index by viewing the three-sphere as an orbi-bundle over the spindle and reducing the three-sphere one-loop determinant, extending a method proposed in [33] to obtain quickly the two-sphere partition function, starting from the three-sphere. In Section 6 we discuss our findings, indicating a number of immediate applications as well as outlining a broader program for continuing the exploration of SQFTs on orbifolds that we initiated here. Further technical material is reported in the appendices.
@@ 2.1 General analysis
which is completely fixed by the Killing spinors $\zeta,{\widetilde{\zeta}}$ via the 1-form $K$ and the scalar $s$. As an example, the covariant Lie derivative $L_{K}$ can be written as follows:
@@
with ${k_{0}}$ being a normalization constant and ${\omega}$ a fugacity for the angular momentum, while $\varphi_{1}=\varphi$ and $\varphi_{2}=\psi$ are angular coordinates with period $2\pi$. We consider a metric that is manifestly invariant with respect to both $\partial_{\varphi}$ and $\partial_{\psi}$:
${\rm d}s^{2}=f^{2}{\rm d}x^{2}+{\mathrm{c}}_{ij}{\rm d}\varphi_{i}{\rm d}\varphi_{j}=\delta_{mn}e^{m}e^{n}~{},$ $i,j=1,2~{},\qquad m,n=1,2,3~{},$ (2.18)
where the metric functions $f=f(x)$ and ${\mathrm{c}}_{ij}={\mathrm{c}}_{ij}(x)$ are allowed to complex values and are independent of $\varphi_{1,2}$. The dreibein $e^{m}$ are
@@
which define a frame regarding the three-dimensional orbifold as a circle fibration over a two-dimensional orbifold parametrized by ${\left({x,\varphi}\right)}$ with line element ${\rm d}s^{2}_{(2)}$, namely
@@
where we parametrize the phase factor $\theta=\theta(x,\varphi,\psi)$ as
@@
$\zeta_{\alpha}={\rm e}^{{\rm i}\theta}\begin{pmatrix}u_{1}\\ -u_{2}\end{pmatrix}_{\alpha}~{},\qquad\widetilde{\zeta}_{\alpha}=-{\rm e}^{-{\rm i}\theta}\begin{pmatrix}u_{2}\\ u_{1}\end{pmatrix}_{\alpha}~{},$ (2.28)
@@ 2.2 $\mathbb{\Sigma}\times S^{1}$ with twist and anti-twist
Hence, an open neighbourhood of $x=+1$ can be parametrized by the complex coordinate $z_{+}$ satisfying
@@
where the identification involving the root of unity $w_{+}$ makes manifest that ${\mathcal{U}_{+}}$ is isomorphic to $\mathbb{C}/{n_{+}}$. On the other hand, near the south pole of the spindle at $x=-1$ we have
$e^{1}_{S}=-\frac{{n_{-}}{\rm d}x}{\sqrt{2{\left({1+x}\right)}}}=-{\rm d}{\rho_{+}}~{},\qquad e^{2}_{S}=\sqrt{2{\left({1+x}\right)}}{\rm d}\varphi=\frac{{\rho_{-}}}{{n_{-}}}{\rm d}\varphi~{},$ ${\rho_{-}}={n_{-}}\sqrt{2{\left({1+x}\right)}}~{}.$ (2.37)
@@
meaning that ${\mathcal{U}_{-}}\cong\mathbb{C}/{n_{-}}$. The minus sign in the exponential of $z_{-}={\rho_{-}}e^{-{\rm i}\varphi/{n_{-}}}$ takes into account the reversed orientation of ${\mathcal{U}_{-}}$ with respect to that of ${\mathcal{U}_{+}}$. A necessary condition for the orbifold $\mathbb{\Sigma}\times S^{1}$ to be supersymmetric is that the magnetic flux of the $U(1)$ R-symmetry field $A^{C}$ through the spindle takes the form
$\frac{1}{2\pi}\int_{\mathbb{\Sigma}}{\rm d}A^{C}=\frac{1}{2}{\left({\frac{1}{{n_{-}}}+\frac{{\mathfrak{s}}}{{n_{+}}}}\right)}=\frac{\chi_{\mathfrak{s}}}{2}~{},$ ${\mathfrak{s}}=\pm 1~{}.$ (2.39)
The values of the sign ${\mathfrak{s}}=\pm 1$ represent two inequivalent ways to preserve supersymmetry on $\mathbb{\Sigma}\times S^{1}$. In the case of ${\mathfrak{s}}=+1$, the magnetic flux of the R-symmetry background field $A^{C}$ equals half of the orbifold Euler characteristic of the spindle $\chi_{\rm orb}={n_{-}}^{-1}+{n_{+}}^{-1}$. This configuration is referred to as topological twist as it is the orbifold generalization of the analogous setup found in spheres [35, 36] and Riemann surfaces [37]. On the other hand, when ${\mathfrak{s}}=-1$, the magnetic flux of the R-symmetry background field equals half of $\chi_{-}={\left({{n_{-}}^{-1}-{n_{+}}^{-1}}\right)}$. This specific configuration is dubbed anti-twist to highlight its contrast with the topological twist, primarily due to the presence of a relative minus sign between ${n_{+}}^{-1}$ and ${n_{-}}^{-1}$, resulting in $\chi_{-}$ being radically distinct from $\chi_{\rm orb}=\chi_{+}$. In the special case where ${n_{+}}={n_{-}}=1$ the spindle become a smooth sphere, and the anti-twisted $\mathbb{\Sigma}\times S^{1}$ configuration reduces to the generalized superconformal index background studied in [38, 39], characterized by the absence of magnetic flux for the R-symmetry connection. Eventually, after shrinking to zero the radius of $S^{1}$ in $\mathbb{\Sigma}\times S^{1}$ and keeping ${n_{+}}={n_{-}}=1$, the anti-twisted spindle becomes the supersymmetric sphere with no R-symmetry twist explored in [33].
The Killing spinors corresponding to the twisted $\mathbb{\Sigma}\times S^{1}$ geometry, occurring if ${\mathfrak{s}}=+1$, are
@@
where $\alpha_{2}|_{N,S}$ is the value of $\alpha_{2}$ in the patch covering $x=\pm 1$, respectively.
Conversely, the Killing spinors associated with the anti-twisted $\mathbb{\Sigma}\times S^{1}$ geometry, which occurs when ${\mathfrak{s}}=-1$, exhibit a formal similarity to the Killing spinors corresponding to the twisted background. However, they differ in terms of their components $u_{1,2}$:
@@
where again $\alpha_{2}|_{N,S}$ are the values of $\alpha_{2}$ in the patches that respectively include the poles $x=\pm 1$.

### v2

@@ Abstract
We consider three-dimensional ${\mathcal{N}}=2$ supersymmetric field theories defined on general complex-valued backgrounds of Euclidean new minimal supergravity admitting two Killing spinors of opposite $R$-charges. We compute partition functions for theories defined on general circle bundles over spindles $\mathbb{\Sigma}$, including $\mathbb{\Sigma}\times S^{1}$ as well as branched and squashed lens spaces, thus obtaining novel observables characterizing three-dimensional supersymmetric gauge theories. We discuss both twisted and anti-twisted theories compactified on $\mathbb{\Sigma}\times S^{1}$ and demonstrate that their partition functions are encoded by a single formula that we refer to as the spindle index, unifying and generalizing superconformal and topologically twisted indices in the limit where orbifold singularities are absent. Furthermore, we test our new index using non-perturbative dualities and obtain one-loop determinants of two-dimensional supersymmetric gauge theories compactified on the spindle.
@@ 1 Introduction
The investigation of supersymmetric quantum field theories (SQFTs) stands as a longstanding pursuit in theoretical physics as it represents an invaluable avenue for exploring non-trivial phenomena, particularly in the regime of strong interactions, where perturbation theory fails. From a physics standpoint the exact computation of SQFTs’ partition functions on curved backgrounds, made possible through localization techniques [1], has produced profound insights on non-perturbative aspects of gauge theories, including the discovery of dualities among seemingly completely different models. From a mathematical point of view, such exact partition functions may serve as generating functions for a number of interesting invariants characterizing the manifolds where the theory is defined, see e.g. [2] and references therein for a recent review.
@@
Partition functions of SQFTs defined on the direct product of a circle and a compact manifold play a special role as they correspond to supersymmetric indices, which may be interpreted in terms of “counting” of various quantities. The study of indices has revealed unforeseen correlations across different areas of physics and mathematics, establishing connections between SQFTs, the geometry of the moduli space in the context of string theory, as well as in purely mathematical fields such as symplectic and algebraic geometry and representation theory, including Gromov-Witten invariants and characters of infinite-dimensional algebras. Recent advancements in localization techniques have yielded significant breakthroughs, facilitating explicit computations of supersymmetric indices as well as other partition functions corresponding to manifolds with diverse topology and dimension.
@@
Supersymmetric partition functions play also an important role in the AdS/CFT correspondence, where they provide precision checks of dualities between gravitational backgrounds and holographically dual field theories. The most striking instance of this aspect is the use of supersymmetric indices to perform a precise counting of the microstates of black holes, thereby offering a glimpse into the fundamental tenets of quantum gravity. In particular, the large-$N$ limit of supersymmetric indices has been extensively used to obtain a microscopic derivation of the entropy of supersymmetric black holes [3, 4, 5, 6]. Several other works have studied different asymptotic limits and determined subleading corrections, matching these to higher-derivative or quantum corrections in supergravity [7, 8, 9, 10]. Another set of precision tests in holography involves matching the large-$N$ limit of partition functions of three-dimensional ${\cal N}=2$ gauge theories on manifolds with $S^{3}$ topology [11, 12, 13] to various supergravity solutions [14, 15, 16, 17, 18].
@@
In the past two years, starting with [19], a substantial body of literature has demonstrated that supergravity solutions displaying orbifold singularities should not be discarded. On the contrary, such solutions are expected to play a crucial role as non-trivial saddle points in the quantum-gravity path integral or in its semi-classical limit. In this extended landscape of solutions supersymmetry can be realized in novel fashions, raising the intriguing question of how typical these supergravity backgrounds are. One of the simplest and perhaps most interesting scenarios is that of four-dimensional supersymmetric [^1] See also [20] for a discussion of supersymmetry of the solutions of [21, 22]. and accelerating black holes [21, 22], wherein the horizon is a spindle $\mathbb{\Sigma}$ [23], with $\mathbb{\Sigma}=\mathbb{WCP}^{1}_{{\left[{{n_{+}},{n_{-}}}\right]}}$ being the two-dimensional weighted complex projective line whose topology is that of a two-sphere with conical singularities $\mathbb{C}/\mathbb{Z}_{n_{\pm}}$ at the poles. The conformal boundary of these geometries comprises a metric on $\mathbb{\Sigma}\times\mathbb{R}_{t}$ together with a gauge field with “anti-twisted” flux on $\mathbb{\Sigma}$; therefore, via holography, their microstates are expected to be captured by the supersymmetric partition function of the dual ${\cal N}=2$ three-dimensional field theories defined on a rigid supersymmetric background on $\mathbb{\Sigma}\times S^{1}$ [24]. In this work we will discuss in detail the calculation of this new observable, that we named spindle index in [25].
@@
More broadly, the recent supergravity constructions involving orbifolds prompted a number of related questions concerning superconformal field theories (SCFTs). For example, it would be interesting to systematically understand the data defining a SCFT at an orbifold point and to identify the additional properties that one can hope to detect on such backgrounds. For SCFTs in even space-time dimension some progress has been made by studying the ’t Hooft anomalies of theories compactified on orbifolds, extending anomaly-polynomial techniques to spaces with conical singularities [19, 26, 27]. This approach is somewhat insensitive to the detailed description of the microscopic fields at the orbifold singularities, suggesting that a similar point of view may be fruitful for studying field theories on orbifolds also in odd space-time dimensions. In particular, we expect that orbifold indices [28] and other orbifold partition functions should unveil novel features of SQFTs in both odd and even space-time dimension, also beyond the context of holography.
Although one of the main motivations of this work was the microstate counting of the supersymmetric and accelerating AdS4 black holes found in [23, 29], in this paper we shall discuss a much more general setting, for which the “anti-twisted” spindle index, relevant for the black holes of [23], is a particular case. Indeed, we will consider ${\mathcal{N}}=2$ supersymmetric field theories placed on compact three-orbifolds and compute their full supersymmetric partition functions. Technically, a key ingredient in our calculations is the equivariant index theorem on orbifolds [30, 31]. We expect that our methods can be readily extended to other setups, including theories in dimension higher than three and possibly incorporating boundaries.
@@
In Section 2 we characterize a very general class of complex, local and rigid supersymmetric backgrounds. We will show that these comprise all the known supersymmetric geometries in three dimensions, as well as several generalizations that we discuss in detail. For instance, we describe general backgrounds on $\mathbb{\Sigma}\times S^{1}$, with both twist and anti-twist for the background $R$-symmetry connection. Moreover, the conformal boundary of the accelerating black hole of [23] as well as its complex deformation [24] are shown to fit into our general canonical form. In Section 3 we calculate the full supersymmetric partition function of three-dimensional ${\cal N}=2$ Chern-Simons-matter theories on $\mathbb{\Sigma}\times S^{1}$, whose one-loop determinants are obtained by employing both the unpaired-eigenvalues method and the equivariant index theorem on orbifolds. The ultimate outcome encompasses anti-twisted and twisted partition functions of gauge theories on $\mathbb{\Sigma}\times S^{1}$, demonstrating that the final expression can be written in a unified form, corresponding to the spindle index [25]. We also include flavour fugacities as well as several classical terms, which were not discussed in [25], and discuss subtleties related to the breaking of gauge invariance and its restoration through the introduction of effective terms. Our final result for the spindle index is tested via some simple field-theory dualities that hold in a form non-trivially affected by the presence of the orbifold singularities of $\mathbb{\Sigma}$. As a further application of our formalism, in Section 4 we compute the full supersymmetric partition function of three-dimensional ${\cal N}=2$ gauge theories on a branched and squashed lens space, again finding a result that unifies and generalises previous constructions regarding non-trivial circle (orbi-)bundles over two-dimensional (orbifold) Riemann surfaces. In particular, our partition functions encompass the path integrals of gauge theories on the squashed lens space [32] and on the branched three-sphere [33] as special cases. Features of this observable and its possible applications will be discussed further in Section 6. In Section 5 we show how to recover the one-loop determinants relevant for the spindle index by viewing the three-sphere as an orbi-bundle over the spindle and subsequently reducing the three-sphere one-loop determinant. This procedure extends a method proposed in [34] to quickly obtain the two-sphere partition function out of the three-sphere one. In Section 6 we summarize and discuss our findings, providing a number of immediate applications and outlining a broader program for continuing the exploration of SQFTs on orbifolds that we initiated here. Further technical material is reported in the appendices.
@@ 2.1 General analysis
which is completely fixed by the Killing spinors $\zeta,{\widetilde{\zeta}}$ via the 1-form $K$ and the scalar ${v}$. As an example, the covariant Lie derivative $L_{K}$ can be written as follows:
@@
with $k_{0}$ being a normalization constant and ${\omega}$ a fugacity for the angular momentum, while $\varphi_{1}=\varphi$ and $\varphi_{2}=\psi$ are angular coordinates with period $2\pi$. In these coordinates, the Killing vector equation $\nabla_{(\mu}K_{\nu)}=0$ reads
[…]
which, if the metric is real, implies that its components are independent of the coordinates $\varphi$, $\psi$. More generally, we require that the metric components are periodic functions of $\varphi$, $\psi$. Expanding in Fourier modes in one of the two coordinates, we have
$g_{\mu\nu}(\varphi,\psi)\,=\,\sum_{n}e^{{\rm i}n\psi}g^{(n)}_{\mu\nu}(\varphi)\,,$ (2.19)
so that
[…]
Thus, $g_{\mu\nu}(\varphi,\psi)$ cannot periodic in $\varphi$ and the only admissible solution is that the metric does not depend on $\varphi$ and $\psi$. Denoting by $x$ the third coordinate, we then have that the metric takes the form
${\rm d}s^{2}=f^{2}{\rm d}x^{2}+{\mathrm{c}}_{ij}({\rm d}\varphi_{i}+c_{i}{\rm d}x)({\rm d}\varphi_{j}+c_{j}{\rm d}x)~{},$ $i,j=1,2~{},$ (2.21)
where the metric functions $f=f(x)$, $c_{i}(x)$, and ${\mathrm{c}}_{ij}={\mathrm{c}}_{ij}(x)$ are allowed to take complex values and are independent of $\varphi_{1,2}$. After reabsorbing $c_{i}(x)$ by a (complex) diffeomorphism, we arrive at the form
${\rm d}s^{2}=f^{2}{\rm d}x^{2}+{\mathrm{c}}_{ij}{\rm d}\varphi_{i}{\rm d}\varphi_{j}~{}.$ (2.22)
Introducing the dreibein $e^{m}$
@@
the three-dimensional orbifold can be viewed as a circle fibration over a two-dimensional orbifold parametrized by ${\left({x,\varphi}\right)}$ with line element ${\rm d}s^{2}_{(2)}$, namely
@@
where we parametrize the phase factor $\theta=\theta(x,\varphi,\psi)$ as [^2] Notice that $\alpha_{2,3}|_{\rm here}=\alpha_{1,2}|_{\rm there}$ with respect to [25].
@@
$\zeta_{\alpha}={\rm e}^{{\rm i}\theta}\begin{pmatrix}u_{1}\\ -u_{2}\end{pmatrix}_{\alpha}~{},\qquad\widetilde{\zeta}_{\alpha}=-{\rm e}^{-{\rm i}\theta}\begin{pmatrix}u_{2}\\ u_{1}\end{pmatrix}_{\alpha}~{},$
@@ 2.2 $\mathbb{\Sigma}\times S^{1}$ with twist and anti-twist
Hence, in an open neighbourhood ${\mathcal{U}_{+}}$ of $x=+1$ the spindle $\mathbb{\Sigma}$ can be parametrized by the complex coordinate $z_{+}$ satisfying
@@
where the identification involving the root of unity $w_{+}$ makes manifest that the topology of ${\mathcal{U}_{+}}$ is that of $\mathbb{C}/\mathbb{Z}_{{n_{+}}}$. On the other hand, near the south pole of the spindle at $x=-1$ we have
$e^{1}_{{\mathcal{U}_{-}}}=-\frac{{n_{-}}{\rm d}x}{\sqrt{2{\left({1+x}\right)}}}=-{\rm d}{\rho_{+}}~{},\qquad e^{2}_{{\mathcal{U}_{-}}}=\sqrt{2{\left({1+x}\right)}}{\rm d}\varphi=\frac{{\rho_{-}}}{{n_{-}}}{\rm d}\varphi~{},$ ${\rho_{-}}={n_{-}}\sqrt{2{\left({1+x}\right)}}~{}.$ (2.36)
@@
meaning that ${\mathcal{U}_{-}}\cong\mathbb{C}/\mathbb{Z}_{{n_{-}}}$. The minus sign in the exponential of $z_{-}={\rho_{-}}e^{-{\rm i}\varphi/{n_{-}}}$ takes into account the reversed orientation of ${\mathcal{U}_{-}}$ with respect to that of ${\mathcal{U}_{+}}$. A necessary condition for the orbifold $\mathbb{\Sigma}\times S^{1}$ to be supersymmetric is that the magnetic flux of the $U(1)$ $R$-symmetry field $A^{C}$ through the spindle takes the form
$\frac{1}{2\pi}\int_{\mathbb{\Sigma}}{\rm d}A^{C}=\frac{1}{2}{\left({\frac{1}{{n_{-}}}+\frac{{\upsigma}}{{n_{+}}}}\right)}=\frac{\chi_{\upsigma}}{2}~{},$ ${\upsigma}=\pm 1~{}.$ (2.38)
The values of the sign ${\upsigma}=\pm 1$ represent two inequivalent ways to preserve supersymmetry on $\mathbb{\Sigma}\times S^{1}$. In the case of ${\upsigma}=+1$, the magnetic flux of the $R$-symmetry background field $A^{C}$ equals half of the orbifold Euler characteristic of the spindle $\chi_{\rm orb}={n_{-}}^{-1}+{n_{+}}^{-1}$. This configuration is referred to as topological twist as it is the orbifold generalization of the analogous setup found in spheres [36, 37] and Riemann surfaces [38]. On the other hand, when ${\upsigma}=-1$, the magnetic flux of the $R$-symmetry background field equals half of $\chi_{-}={\left({{n_{-}}^{-1}-{n_{+}}^{-1}}\right)}$. This specific configuration is dubbed anti-twist to highlight its contrast with the topological twist, primarily due to the presence of a relative minus sign between ${n_{+}}^{-1}$ and ${n_{-}}^{-1}$, resulting in $\chi_{-}$ being radically distinct from $\chi_{\rm orb}=\chi_{+}$. In the special case where ${n_{+}}={n_{-}}=1$ the spindle become a smooth sphere, and the anti-twisted $\mathbb{\Sigma}\times S^{1}$ configuration reduces to the generalized superconformal index background studied in [39, 40], characterized by the absence of magnetic flux for the $R$-symmetry connection. Eventually, after shrinking to zero the radius of $S^{1}$ in $\mathbb{\Sigma}\times S^{1}$ and keeping ${n_{+}}={n_{-}}=1$, the anti-twisted spindle becomes the supersymmetric sphere with no $R$-symmetry twist explored in [34].
The Killing spinors corresponding to the twisted $\mathbb{\Sigma}\times S^{1}$ geometry, occurring if ${\upsigma}=+1$, are
$\zeta_{\alpha}=e^{{\rm i}\theta}\begin{pmatrix}u_{1}\\ -u_{2}\end{pmatrix}_{\alpha}~{},\qquad\widetilde{\zeta}_{\alpha}=-e^{-{\rm i}\theta}\begin{pmatrix}u_{2}\\ u_{1}\end{pmatrix}_{\alpha}~{},$
@@
where $\alpha_{2}|_{{\mathcal{U}_{+}},{\mathcal{U}_{-}}}$ is the value of $\alpha_{2}$ in the patch covering $x=\pm 1$, respectively.
Conversely, the Killing spinors associated with the anti-twisted $\mathbb{\Sigma}\times S^{1}$ geometry, which occurs when ${\upsigma}=-1$, exhibit a formal similarity to the Killing spinors corresponding to the twisted background. However, they differ in terms of their components $u_{1,2}$:
@@
where again $\alpha_{2}|_{{\mathcal{U}_{+}},{\mathcal{U}_{-}}}$ are the values of $\alpha_{2}$ in the patches that respectively include the poles $x=\pm 1$.